A spherical liquid drop of radius R is divided into eight equql droplets If surface tension is T then work done in the process will be

Grade:12

2 Answers

askiitiansexpert soumyajit_iitkanpur
14 years ago

Volume of bigger drop = Total volume of 8 droplets

(4/3)*pi* R3   =  8 (4/3)* pi* r3

=> r = R/2

Change in surface area in this process = ds = 4* pi * [8*r2 - R2] = 4* pi* R2

Hence work done = T.ds = T* 4*pi*R2

 

Soumyajit Mondal
5 years ago
V=8v
R³=n*r³
Work done =change in surface energy
ΔU=T * Δ Α
   = T * 4 π ( n r² - R²)
   = T* 4π  n ( R² * n^ (-2/3) - R²)
   = T 4π R² ( n^{1/3} - 1 )